Zoom stock has a beta of 1.46. The risk-free rate of return is 3.07 percent and the market rate of return is 11.81 percent. What is the amount of the risk premium on Zoom stock?
A) 8.09%
B) 12.76%
C) 9.59%
D) 10.25%
E) 17.24%
Answer #1
B) 12.76%
Beta, B 1.46
Return on Market, Rm 11.81%
Risk Free Rate, Rf 3.07%
Risk Premium = Beta(Rm-Rf)
Risk Premium = 1.46(11.81% - 3.07%)
Risk Premium = 1.46 * 8.74%
Risk Premium = 12.76%